OpenStudy (anonymous):

2a+3/5= a/2 then a+3= ?

OpenStudy (blank ):

Which one is it?: \[2a + \frac{ 3 }{5 }=\frac{ a }{ 2 }\]or \[\frac{ 2a+3 }{ 5 } = \frac{ a }{2 }\]

OpenStudy (anonymous):

multiply with 10 on both sides and solve it